OpenStudy (anonymous):

1. What is the value of n in 12/(n+2)=7/(n-3)? a. 1 b. 5 c. 10 d. 20 2. What is the value of k in (2k+4)/10=(k-2)/3 a. 1 1/3 b. 2 c. 6 d. 8 3. What values of x makes this equation true? x/(2x-5)=2/(x-3) a. x=2 and x=5 b. x=2 and x= -5 c. x= -2 and x=5 d. x= -2 and x= -5 4. Solve the equation below for a. 4/(a+9) + 2/a = 15/5a a. 3 b. 9 c. 15 d. 454

OpenStudy (anonymous):

5. Solve the equation below for x. x/(16-x)=2/(x+2) a. x=2 and x=8 b. x=4 and x= -8 c. x=8 and x= -4 d. x= 8 and x=2 6. Solve the equation below for x. 24/(x-3) - 30/(x+3) = 120/(x+3)(x-3) a. 3 b. 4 c. 5 d. 7 7.Solve the equation below for b. (2-b)/30 = (17+2b)/45 a. 6 b. 4 c. -4 d. -6 8. Solve the equation below for x. (12-2x)/x = (x+1)/2 a. x=6 and x= -4 b. x=3 and x= -8 c. x= -3 and x=8 d. x= -6 and x=4

OpenStudy (anonymous):

For the first 2, cross multiply then solve for the variable.